A 500-turn solenoid have a cross-sectional area 20 m^2. If its self-inductance



is 20H, then find the length of the solenoid.

1
Expert's answer
2022-04-28T16:41:53-0400
"L=\frac{\mu_0N^2A}{l}"

"l=\frac{\mu_0N^2A}{L}=\frac{4\pi*10^{-7}*500^2*20}{20}=0.314\:\rm m"